# Fractional Thouless pumping of solitons: a unique manifestation of bulk-edge correspondence of nonlinear eigenvalue problems

### v2

@@ Abstract
Recent foundational studies have established the bulk-edge correspondence for nonlinear eigenvalue problems using auxiliary eigenvalues $\hat{H}\Psi=\omega S(\omega)\Psi$, spanning both linear [ T. Isobe et al., Phys. Rev. Lett. 132, 126601 (2024)] and nonlinear [ Chenxi Bai and Zhaoxin Liang, Phys. Rev. A. 111, 042201 (2025)] Hamiltionians. This raises a fundamental question: Can eigenvalue nonlinearity generate observable physical phenomena absent in conventional approaches ($\hat{H}\Psi=E\Psi$)? In this work, we demonstrate the first uniquely nonlinear manifestation of this correspondence: fractional Thouless pumping of solitons. Through systematic investigation of nonlinear Thouless pumping in an extended Rice-Mele model with next-nearest-neighbor (NNN) couplings, we discover that the NNN parameters can induce fractional topological phases despite quantized topological invariants from conventional approaches. Crucially, these fractional phases are explained within the auxiliary eigenvalue framework, linking nonlinear spectral features directly to bulk-boundary correspondence. This work reveals novel phenomena emerging from the interplay between nonlinearity and NNN couplings, providing key insights for designing topological insulators and controlling quantum edge states in nonlinear regimes.
@@ I Introduction
Conceptually going beyond above conventional approaches, a recent breakthrough [63] has established a bridge between bulk-edge correspondence and eigenvalue nonlinearity in linear Hamiltonians throuth auxiliary eigenvalues $\hat{H}\Psi=\omega S(\omega)\Psi$. This framework demonstrates how topological edge states of auxiliary eigenstates emerge as physical edge states under weak but finite nonlinearity (requiring monotonic $\omega$-dependence), rigorously extending bulk-edge correspondence to nonlinear eigenvalue constraints [64, 65, 66, 67, 68, 69, 70, 71, 72]. Building on this foundation, our work [73] generalizes the auxiliary eigenvalue approach to intrinsic nonlinear Hamiltonians, revealing an anomalous bulk-edge correspondence rooted in these nonlinear eigenvalues. Collectively, Refs. [63, 73] provide a unified description of bulk-edge physics across linear and nonlinear systems via nonlinear eigenvalue problems. This raises a key question: Can this framework predict observable phenomena inaccessible to conventional linear bulk-edge paradigms?
@@
Figure 1: (a) Schematic diagram of the nonlinear extended RM model. The dotted box indicates the unit cell composed of sublattices A and B. (b) Chern number $\mathcal{C}$ topological phase diagram of the linear extended RM Hamiltonian (1) with $g=0$ under $H_{\text{ERM}}\Psi=E\Psi$. (c) Chern number $\mathcal{C}$ topological phase diagram of the linear extended RM Hamiltonian (1) with $g=0$ under $H_{\text{ERM}}\Psi=\omega S(\omega)\Psi$. The red dot represent the fractional Chern number of $\mathcal{C}=-\frac{1}{2}$ in specific NNN coupling parameter regimes of $t_{a}$ and $t_{b}$. (b,c) The parameters are given as $J=1$, $\delta=0.5$, $\Delta=1$, $T=2000\pi$, and $\omega=10^{-3}$.
The organizational structure of this paper is outlined as follows: In Sec. II, we provide an in-depth introduction to the Hamiltonian and Schrödinger equation of the nonlinear extended RM model. In Sec. III, we investigate the energy spectra and phase diagrams of the nonlinear extended RM model under both linear and nonlinear eigenvalue conditions, examining the effects of nonlinearity strength and NNN parameters on these properties. Sec. IV discusses bulk-edge correspondence in the anomalous eigenvalue’s nonlinearity regime, elucidating the role of model parameters. Finally, in Sec. V, we summarize the results of our investigation.
@@ II Nonlinear extended Rice-Mele model
[…]
-- $\Delta\cos\left(\omega t\right)\sum_{n=1}^{N}\left(|\Psi_{n,A}|^{2}-|\Psi_{n,B}|^{2}\right)$
@@
[…]
In above Eqs. (2) and (3), each unit cell consists of two sublattice sites labelled by A and B, and the collective wavefunction $\Psi_{n}=[\psi_{n,A},\psi_{n,B}]$ encapsulates the amplitude on the $n$-th dimer. The intracell and intercell couplings are $J_{1}=-J-\delta\sin(\omega t)$ and $J_{2}=-J+\delta\sin(\omega t)$ respectively, where $J$ sets the uniform hopping amplitude. The staggered sublattice potential $\Delta\sin(\omega t)$ introduces an energy offset between sites, which is a periodically modulates dimerization pattern, while nonlinear interactions are introduced via a Kerr-type term with strength $g$, favoring self-localization [76]. The NNN Hamiltonian $H_{\text{NNN}}$ contains hopping amplitudes $t_{a}$ and $t_{b}$ along sublattices A and B. Crucially, the model reduces to the standard nonlinear Rice-Mele Hamiltonian [76] when $t_{a}=t_{b}=0$. Finally, we note that Hamiltonian (2) provides a direct theoretical framework for recent experiments [56] investigating interaction-induced Thouless pumping in dynamical superlattices or pulsed light propagation in waveguide arrays [76].
@@ II.1 Topological phase diagram for linear and nonlinear eigenvalue problems
[…]
@@
Next, there are two kinds way of defining the eigenvalue problems of Hamiltonian (1) [63, 73], i. e. one is based on the conventional approaches of $H_{\text{ERM}}\Psi=E\Psi$; The other is referred as to the the bulk-edge correspondence of nonlinear eigenvalue problems of linear Hamiltonian (1) with $g=0$ defined as [63, 73]
$H_{\text{ERM}}(k,t)\Psi=\omega S(\omega,k)\Psi.$ (4)
@@
We begin by defining the auxiliary matrix $P\left(\omega,k\right)=H_{\text{ERM}}\left(\omega,k\right)-\omega S\left(\omega,k\right)$, whose kernel, $P\left(\omega,\mathbf{k}\right)\varPsi=0$, reproduces the nonlinear eigen-problem stated in Eq. (4). To elucidate the bulk-edge correspondence encoded in this equation, we promote the equality to an eigen-problem $P\left(\omega,k\right)\varPsi=\lambda\varPsi$, with $\lambda\in\mathbb{R}$ serving as a bookkeeping parameter. Physical relevance is restricted to zero-eigenvalue subspace $\lambda=0$, so the central task reduces to solving $P\left(\omega,k\right)\varPsi=\lambda\varPsi$ at $\lambda=0$. The overlap matrix $S$ entering Eq. (4) is constructed as $S=I-M_{S}\sigma_{z}$, where $\sigma_{z}$ donates the Pauli z-matrix and $M_{S}\left(\omega\right)=M_{1}\tanh\left(\omega t\right)/\omega$ encodes the nonlinear $\omega$-dependence.
@@
Finally, in Figs. 1(b) and (c), we present the topological phase diagrams depicting the Chern number $\mathcal{C}$ for the linear eigenvalue problem ($H_{\text{ERM}}\Psi=E\Psi$) and the nonlinear eigenvalue problem (Eq. 4), respectively, both evaluated at $g=0$ for $H_{\text{ERM}}$. Within the linear framework ($H_{\text{ERM}}\Psi=E\Psi$), the Chern number $\mathcal{C}$ remains quantized at $1$ (Fig. 1(b)). In contrast, the nonlinear eigenvalue problem $H_{\text{ERM}}\Psi=\omega S(\omega)\Psi$ yields fractional $\mathcal{C}$, exemplified by $\mathcal{C}=-\frac{1}{2}$ at $t_{a}=0.5$, $t_{b}=-0.5$ (red dot, Fig. 1(c)). Our subsequent goal is to demonstrate how the bulk-edge correspondence, modified by the eigenvalue nonlinearity evident in Fig. 1(c), gives rise to the observable physical phenomenon of fractional soliton pumping. This phenomenon is absent in the linear case shown in Fig. 1(b).
@@ II.2 Nonlinear extended Schrödinger equation
To study soliton dynamics, we derive the equations of motion for the nonlinear extended RM Hamiltonian $H_{\text{ERM}}$ by applying the variational principle $i\partial\Psi_{j}/\partial t=\delta H/\delta\Psi^{*}_{j}$, yielding the following discrete nonlinear Schrödinger equations for sites $j=1,...N-1,N$,
$i\frac{\partial\Psi_{j,A}}{\partial t}=-\left(J+\delta\sin\omega t\right)\Psi_{j,B}-\left(J-\delta\sin\omega t\right)\Psi_{j-1,B}$
$-\left[\Delta\cos\omega t+g\left|\Psi_{j,A}\right|^{2}\right]\Psi_{j,A}\!+\!t_{a}(\Psi_{j-1,A}\!+\!\Psi_{j+1,A}),$ (5)
$i\frac{\partial\Psi_{j,B}}{\partial t}=-\left(J+\delta\sin\omega t\right)\Psi_{j,A}-\left(J-\delta\sin\omega t\right)\Psi_{j-1,A}$
$+\left[\Delta\cos\omega t-g\left|\Psi_{j,B}\right|^{2}\right]\Psi_{j,B}\!+\!t_{b}(\Psi_{j-1,B}\!+\!\Psi_{j+1,B}).$ (6)
@@
Figure 2: (a1)-(d1): Band structures of the nonlinear eigenvalues for the nonlinear extended RM Hamiltonian (1), (a2)-(d2): The position expectation value $\langle X\rangle$ of the soliton as a function of time over a single period. The parameters are fixed as $J=1$, $\delta=0.5$, $\Delta=1$, $t_{a}=0.4$, $t_{b}=-0.1$, $T=2000\pi$, and $\omega=10^{-3}$. (a1)-(d1) [(a2)-(d2)]: The interaction strengths are set to $g=0$, $g=1$, $g=5$ and $g=10$, respectively.
Figure 3: NNN tunneling control of topological transport: (a,c) Nonlinear eigenvalue spectra ($\mu$); (b,d,e) Quantized soliton dynamics. Fixed parameters: $J=1$, $\delta=0.5$, $\Delta=1$, $g=1$, $T=2000\pi$, $\omega=10^{-3}$. Tunneling configurations: (a,b) $t_{a}=0.4$, $t_{b}=-0.1$; (c,d) $t_{a}=0.5$, $t_{b}=-0.5$.
@@ III.1 Nonlinear eigenvalue of the nonlinear Hamiltonian
In this subsection III.1, we are interested into both the quantized nonlinear Thouless pumping and fractional nonlinear Thouless pumping predicted by the topological phase diagram in Figs. 1(b) and (c). This is achieved by numerically solving Eqs. (5) and (6) under a continuous change of the nonlinearity parameter $g$ in Eq. (1) from zero to a finite value. All simulations employ a $100$-unit cell lattice ($200$ sites). The initial soliton state, localized at site $n=0$, is prepared using the iterative self-consistent algorithm from Refs. [73, 76] with trial wavefunction $\Psi_{0}=\text{sech}\left[\lvert x-100\rvert/5\right]$. Here, we use two distinct methodologies to solve Eqs. (5) and (6) numerically. In the first approach entitled the method of dynamically evolved soltions, we use a fourth-order Runge-Kutta algorithm to numerically propagate the initial soliton profile through time evolution governed by Eqs. (5) and (6). This method directly simulates the dynamical evolution of the wavefunction under physical time-dependent driving. In the second approach called by the method of instantaneous solitons, an iterative self-consistent scheme is employed to converge to the steady-state solution at each time slice $t$, bypassing explicit time propagation. In the adiabatic regime ($\omega\rightarrow 0$), the two approaches converge to the same soliton trajectory, confirming numerical consistency.
@@
The quantized soliton transport in Fig. 2 originates from the Chern number (see Fig. 1(b)) in the noninteracting limit ($g=0$) [54, 73, 76, 84]. Specifically, under weak nonlinearity (Fig. 2(b1)), the soliton occupies the lowest band of Hamiltonian (1), whose linear dynamical Chern number equals $1$ (Fig. 1(b)). Consequently, the bulk-edge correspondence for the nonlinear extended RM Hamiltonian (1) yields quantized displacement of two lattice sites per cycle (Fig. 2(b2)). In contrast, strong nonlinearity (Fig. 2(d2)) induces Rabi oscillations between the two lowest bands. The vanishing sum of dynamical Chern numbers ($\mathcal{C}_{1}+\mathcal{C}_{2}=0$) then freezes soliton transport. For moderate interactions $g$ (Fig. 2(c2)), transport remains quantized when calculated via the instantaneous soliton method. However, dynamically evolved soliton trajectories exhibit position jumps (Fig. 2(c2)) due to emergent self-intersecting bands in the intermediate nonlinear regime. This band topology, characterized by loop structures [85], is visualized in Fig. 2(c1).
We next examine how NNN tunneling amplitudes $t_{a}$ and $t_{b}$ in Hamiltonian (1) influence quantized soliton transport, maintaining fixed weak nonlinearity ($g=1$). These effects are systematically investigated through Fig. 3. We first analyze the case $t_{a}=0.4$, $t_{b}=-0.1$. As demonstrated in Fig. 3(a), gapless edge states persist due to preserved inversion symmetry, resulting in stable soliton oscillations (Fig. 3(b)). Crucially, the ground-state soliton exhibits a pumping value of $2$ at this parameter point—consistent with Chern number $\mathcal{C}=1$ (corresponding to two atoms per unit cell)—and displays perfect adiabatic following, where dynamically evolved solitons precisely match their instantaneous counterparts. When the absolute values of $t_{b}$ becomes comparable to $t_{a}$ ($t_{a}=0.5$, $t_{b}=-0.5$), inversion symmetry breaks. This symmetry breaking opens a bandgap in the $\mu$-spectrum (Fig. 3(c)) and hybridizes edge states into bulk modes, thereby disrupting the adiabatic pathway as evidenced in Fig. 3(d). By selecting distinct NNN tunneling amplitudes $\{t_{a},t_{b}\}$ and superposing their corresponding soliton trajectories as shown in Fig. 3(e), we observe that all pumping pathways exhibit perfect adiabatic following except along the symmetry-broken line $t_{a}=-t_{b}=0.5$. Crucially, adiabatic breakdown occurs exclusively on this $t_{a}=-t_{b}$ trajectory. This reveals how NNN tunneling parameters reconfigure the system’s topological phase diagram.
Figure 4: NNN tunneling modulation of topological phases: (a,c) Nonlinear eigenvalue spectra ($\mu$); (b,d,e) Quantized soliton transport dynamics. Fixed parameters: $J=1$, $\delta=0.5$, $\Delta=1$, $g=6.215$, $T=2000\pi$, $\omega=10^{-3}$. Parameter configurations: (a,b) $t_{a}=0.4$, $t_{b}=-0.1$; (c,d) $t_{a}=0.5$, $t_{b}=-0.5$.

### v3

@@ Abstract
Recent foundational studies have established the bulk-edge correspondence for nonlinear eigenvalue problems using auxiliary eigenvalues $\hat{H}\Psi=\omega S(\omega)\Psi$, spanning both linear [ T. Isobe et al., Phys. Rev. Lett. 132, 126601 (2024)] and nonlinear [ Chenxi Bai and Zhaoxin Liang, Phys. Rev. A. 111, 042201 (2025)] Hamiltionians. This progress prompts a fundamental question: Can eigenvalue nonlinearity generate observable physical phenomena absent in conventional approaches ($\hat{H}\Psi=E\Psi$)? In this work, we address this question by demonstrating the first uniquely nonlinear manifestation of the bulk-edge correspondence: fractional Thouless pumping of solitons. Through systematic investigation of nonlinear Thouless pumping in an extended Rice-Mele model incorporating next-nearest-neighbor (NNN) couplings, we uncover that NNN interaction parameters can induce fractional topological phases—even in the presence of quantized topological invariants as predicted by conventional linear approaches. Crucially, these fractional phases are naturally explained within the auxiliary eigenvalue framework, directly linking nonlinear spectral characteristics to the bulk-boundary correspondence. Our findings reveal novel emergent phenomena arising from the interplay between nonlinearity and NNN couplings, providing key insights for the design of topological insulators and the controlled manipulation of quantum edge states in nonlinear regimes.
@@ I Introduction
Conceptually going beyond above conventional approaches, a recent breakthrough [63] has forged a connection between bulk-edge correspondence and eigenvalue nonlinearity in linear Hamiltonians via the auxiliary eigenvalue formulation $\hat{H}\Psi=\omega S(\omega)\Psi$. This framework elucidates how topological edge states of auxiliary eigenstates manifest as physical edge states under weak yet finite nonlinearity (requiring monotonic $\omega$-dependence), rigorously extending the bulk-edge correspondence to nonlinear eigenvalue problems [64, 65, 66, 67, 68, 69, 70, 71, 72]. Building upon this foundation, our work [73] further generalizes the auxiliary eigenvalue approach to intrinsic nonlinear Hamiltonians, revealing an anomalous bulk-edge correspondence inherently tied to these nonlinear eigenvalues. Collectively, Refs. [63, 73] establish a unified description of bulk-edge physics across linear and nonlinear systems through the lens of nonlinear eigenvalue problems. This raises a pivotal question: Can this framework predict observable phenomena inaccessible to conventional linear bulk-edge paradigms?
@@
Figure 1: (a) Schematic diagram of the nonlinear extended RM model. The dotted box indicates the unit cell composed of sublattices A and B. (b) Chern number $\mathcal{C}$ topological phase diagram of the linear extended RM Hamiltonian (1) with $g=0$ under $H_{\text{ERM}}\Psi=E\Psi$. (c) Chern number $\mathcal{C}$ topological phase diagram of the linear extended RM Hamiltonian (1) with $g=0$ under $H_{\text{ERM}}\Psi=\omega S(\omega)\Psi$. The red dot represent the fractional Chern number of $\mathcal{C}=-\frac{1}{2}$ in specific NNN coupling parameter regimes of $t_{a}$ and $t_{b}$. (b,c) The parameters are given as $J=1$, $\delta=0.5$, $\Delta=1$, $T=2000\pi$, and $\omega_{\text{d}}=10^{-3}$
This paper is structured as follows: In Section II, we present a comprehensive introduction to the Hamiltonian formulation and Schrödinger equation governing the nonlinear extended RM model. Section III analyzes the energy spectra and phase diagrams of this model under both linear and nonlinear eigenvalue conditions, with a focus on how nonlinearity strength and NNN coupling parameters influence these properties. In Section IV, we explore the bulk-edge correspondence within the nonlinear eigenvalue regime of anomalous states, elucidating the role of key model parameters in mediating topological phenomena. Finally, Section V summarizes our principal findings and their implications for future research.
@@ II Nonlinear extended Rice-Mele model
$H_{\text{RM}}$ $=$ $\sum_{n=1}^{N}\left(J_{1}\Psi_{n,A}^{*}\Psi_{n,B}+J_{2}\Psi_{n,B}^{*}\Psi_{n+1,A}+\text{H.c.}\right)$ (2)
- $\Delta\cos\left(\omega_{\text{d}}t\right)\sum_{n=1}^{N}\left(|\Psi_{n,A}|^{2}-|\Psi_{n,B}|^{2}\right)$
@@
$\!\!\!\!H_{\text{NNN}}\!\!\!$ $=$ $\!\!\sum_{n=1}^{N-1}\left(t_{a}\Psi_{n,A}^{*}\Psi_{n+1,A}+t_{b}\Psi_{n,B}^{*}\Psi_{n+1,B}\!+\!\text{H.c.}\right).$ (3)
In above Eqs. (2) and (3), each unit cell consists of two sublattice sites labelled by A and B, and the collective wavefunction $\Psi_{n}=[\psi_{n,A},\psi_{n,B}]$ encapsulates the amplitude on the $n$-th dimer. The intracell and intercell couplings are $J_{1}=-J-\delta\sin(\omega_{\text{d}}t)$ and $J_{2}=-J+\delta\sin(\omega_{\text{d}}t)$ respectively, where $J$ sets the uniform hopping amplitude. The staggered sublattice potential $\Delta\cos(\omega_{\text{d}}t)$ introduces an energy offset between sites, which is a periodically modulates dimerization pattern, while nonlinear interactions are introduced via a Kerr-type term with strength $g$, favoring self-localization [76]. The NNN Hamiltonian $H_{\text{NNN}}$ contains hopping amplitudes $t_{a}$ and $t_{b}$ along sublattices A and B. Crucially, the model reduces to the standard nonlinear Rice-Mele Hamiltonian [76] when $t_{a}=t_{b}=0$. Finally, we note that Hamiltonian (2) provides a direct theoretical framework for recent experiments [56] investigating interaction-induced Thouless pumping in dynamical superlattices or pulsed light propagation in waveguide arrays [76].
@@ II.1 Topological phase diagram for linear and nonlinear eigenvalue problems
We first compute the topological phase diagram for the linear extended Rice-Mele model [Hamiltonian (1) with $g=0$]. Under periodic boundary conditions, the momentum-space wavefunctions for each sublattice are obtained by discrete Fourier transformation: $|\Psi_{A}(k,t)\rangle=1/\sqrt{N}\sum_{j=1}^{N}e^{ikj}|\Psi_{j,A}(t)\rangle$ and $|\Psi_{B}(k,t)\rangle=1/\sqrt{N}\sum_{j=1}^{N}e^{ikj}|\Psi_{j,B}(t)\rangle$, with $k=2\pi m/N$ ($m=0,1,\dots,N-1$). The total wavefunction is expressed as a two-component spinor $|\Psi(k,t)\rangle=(|\Psi_{A}(k,t)\rangle,|\Psi_{B}(k,t)\rangle)^{T}$, and then the Hamiltonian $H_{\text{ERM}}(k,t)$ can be written in the form of $H_{\text{ERM}}(k,t)=\langle\Psi(k,t)|h(k,t)|\Psi(k,t)\rangle$. Here, the $h(k,t)$ can expressed in the form of $h(k,t)=h_{\text{I}}I+d(k,t)\cdot\sigma$ with I being the identity matrix, $\sigma=\left(\sigma_{x},\sigma_{y},\sigma_{z}\right)$ representing the Pauli matices, and $h_{\text{I}}=(t_{a}+t_{b})\cos k$, $d_{x}=(J_{1}+J_{2}\cos k)$, $d_{y}=J_{2}\sin k$, and $d_{z}=\left(t_{a}-t_{b}\right)\cos k-\Delta\cos(\omega_{\text{d}}t)$. This Hamiltonian of $h(k,t)$ has broken chiral symmetry and inversion symmetry, and exhibit no other protecting symmetry. The robust quantized pumping arise from the nontrivial Chern number associated with band structures in the combined time-momentum space. The Chern number, defined over the 2D torus $(k,\theta)$ with $\theta=\omega_{\text{d}}t\in[0,2\pi]$, does not rely on symmetry and serves as a topological invariant that ensures the quantization of the pumped charge per cycle, as long as the energy gap remains open during the adiabatic evolution. In more details, the Chern number for the first non-degenerate band is defined as the integral of the Berry curvature over the Brillouin zone [26, 63]: $\mathcal{C}=\frac{1}{2\pi}\int_{0}^{2\pi}dk\int_{0}^{T}dt\left[\nabla_{\vec{R}}\times\mathcal{A}(\vec{R})\right]$, where $\vec{R}=[0,2\pi]\times[0,T]$ and $\mathcal{A}(\vec{R})=i\langle\Psi(k,t)|\nabla_{k}|\Psi(k,t)\rangle$ is the Berry connection.
@@
Next, there are two approaches to defining the eigenvalue problem for the Hamiltonian (1) [63, 73], i. e. one is based on the conventional approaches of $H_{\text{ERM}}\Psi=E\Psi$; The other is referred as to the the bulk-edge correspondence of nonlinear eigenvalue problems of linear Hamiltonian (1) with $g=0$ defined as [63, 73]
$H_{\text{ERM}}(k,t)\Psi=\omega S(\omega)\Psi.$ (4)
@@
We begin by defining the auxiliary matrix $P\left(\omega,k\right)=H_{\text{ERM}}\left(k,t\right)-\omega S\left(\omega\right)$, whose kernel, $P\left(\omega,\mathbf{k}\right)\varPsi=0$, reproduces the nonlinear eigen-problem stated in Eq. (4). To elucidate the bulk-edge correspondence encoded in this equation, we promote the equality to an eigen-problem $P\left(\omega,k\right)\varPsi=\lambda\varPsi$, with $\lambda\in\mathbb{R}$ serving as a bookkeeping parameter. Physical relevance is restricted to zero-eigenvalue subspace $\lambda=0$, so the central task reduces to solving $P\left(\omega,k\right)\varPsi=\lambda\varPsi$ at $\lambda=0$. The overlap matrix $S(\omega)$ entering Eq. (4) is constructed as $S(\omega)=I-M_{S}\sigma_{z}$, where $\sigma_{z}$ donates the Pauli z-matrix and $M_{S}\left(\omega\right)=M_{1}\tanh\left(\omega t\right)/\omega$ encodes the nonlinear $\omega$-dependence.
@@
Finally, in Figs. 1(b) and (c), we present the topological phase diagrams depicting the Chern number $\mathcal{C}$ for the linear eigenvalue problem ($H_{\text{ERM}}\Psi=E\Psi$) and the nonlinear eigenvalue problem (Eq. 4), respectively, both evaluated at $g=0$ for $H_{\text{ERM}}$. Within the linear framework ($H_{\text{ERM}}\Psi=E\Psi$), the value of the Chern number $\mathcal{C}$ remains quantized at 1-1 (Fig. 1(b)). In contrast, the nonlinear eigenvalue problem $H_{\text{ERM}}\Psi=\omega S(\omega)\Psi$ yields fractional $\mathcal{C}$, exemplified by $\mathcal{C}=-\frac{1}{2}$ at $t_{a}=0.5$, $t_{b}=-0.5$ (red dot, Fig. 1(c)). Our subsequent goal is to demonstrate how the bulk-edge correspondence, modified by the eigenvalue nonlinearity evident in Fig. 1(c), gives rise to the observable physical phenomenon of fractional soliton pumping. This phenomenon is absent in the linear case shown in Fig. 1(b).
@@ II.2 Nonlinear extended Schrödinger equation
To study soliton dynamics, we derive the equations of motion for the nonlinear extended RM Hamiltonian $H_{\text{ERM}}$ by applying the variational principle $i\partial\Psi_{j}/\partial t=\delta H/\delta\Psi^{*}_{j}$, yielding the following discrete nonlinear Schrödinger equations for sites j=1,N1,Nj=1,...N-1,N,
$i\frac{\partial\Psi_{j,A}}{\partial t}=-\left(J+\delta\sin\omega_{\text{d}}t\right)\Psi_{j,B}-\left(J-\delta\sin\omega_{\text{d}}t\right)\Psi_{j-1,B}$
$-\left[\Delta\cos\omega_{\text{d}}t+g\left|\Psi_{j,A}\right|^{2}\right]\Psi_{j,A}\!+\!t_{a}(\Psi_{j-1,A}\!+\!\Psi_{j+1,A}),$ (5)
$i\frac{\partial\Psi_{j,B}}{\partial t}=-\left(J+\delta\sin\omega_{\text{d}}t\right)\Psi_{j,A}-\left(J-\delta\sin\omega_{\text{d}}t\right)\Psi_{j-1,A}$
$+\left[\Delta\cos\omega_{\text{d}}t-g\left|\Psi_{j,B}\right|^{2}\right]\Psi_{j,B}\!+\!t_{b}(\Psi_{j-1,B}\!+\!\Psi_{j+1,B}).$ (6)
@@
Figure 2: (a1)-(d1): Band structures of the nonlinear eigenvalues for the nonlinear extended RM Hamiltonian (1), (a2)-(d2): The position expectation value $\langle X\rangle$ of the soliton as a function of time over a single period. The parameters are fixed as $J=1$, $\delta=0.5$, $\Delta=1$, $t_{a}=0.4$, $t_{b}=-0.1$, $T=2000\pi$, and $\omega_{\text{d}}=10^{-3}$. (a1)-(d1) [(a2)-(d2)]: The interaction strengths are set to $g=0$, $g=1$, $g=5$ and $g=10$, respectively.
Figure 3: NNN tunneling control of topological transport: (a,c) Nonlinear eigenvalue spectra ($\mu$); (b,d,e) Quantized soliton dynamics. Fixed parameters: $J=1$, $\delta=0.5$, $\Delta=1$, $g=1$, $T=2000\pi$, $\omega_{\text{d}}=10^{-3}$. Tunneling configurations: (a,b) $t_{a}=0.4$, $t_{b}=-0.1$; (c,d) $t_{a}=0.5$, $t_{b}=-0.5$.
@@ III.1 Nonlinear eigenvalue of the nonlinear Hamiltonian
In this subsection III.1, we are interested into both the quantized nonlinear Thouless pumping and fractional nonlinear Thouless pumping predicted by the topological phase diagram in Figs. 1(b) and (c). This is achieved by numerically solving Eqs. (5) and (6) under a continuous change of the nonlinearity parameter $g$ in Eq. (1) from zero to a finite value. All simulations employ a $100$-unit cell lattice ($200$ sites). To ensure numerical reliability, we have systematically verified through size-convergence studies that all simulations in this work demonstrate the system-size independence. The initial soliton state, localized at site $n=0$, is prepared using the iterative self-consistent algorithm from Refs. [73, 76] with trial wavefunction $\Psi_{0}=\text{sech}\left[\lvert x-100\rvert/5\right]$. Here, we use two distinct methodologies [76, 73] to solve Eqs. (5) and (6) numerically, with detailed methodologies presented in Appendix A. In brief, the first method is called by dynamically evolved soliton method corresponding to direct time integration of Eqs. (5) and (6) using a fourth-order Runge-Kutta algorithm with initial soliton conditions; while, the second method is referred to as instantaneous soliton solution method corresponding to the iterative determination of stationary states through successive approximation, using the previous time-step solution as the initial trial wavefunction.
@@
The quantized soliton transport in Fig. 2 originates from the Chern number (see Fig. 1(b)) in the noninteracting limit ($g=0$) [54, 73, 76, 84]. Specifically, under weak nonlinearity (Fig. 2(b1)), the soliton occupies the lowest band of Hamiltonian (1), whose linear dynamical Chern number equals 1-1 (Fig. 1(b)). Consequently, the bulk-edge correspondence for the nonlinear extended RM Hamiltonian (1) yields quantized displacement of two lattice sites per cycle (Fig. 2(b2)). In contrast, strong nonlinearity (Fig. 2(d2)) induces Rabi oscillations between the two lowest bands. The vanishing sum of dynamical Chern numbers ($\mathcal{C}_{1}+\mathcal{C}_{2}=0$) then freezes soliton transport. For moderate interactions $g$ (Fig. 2(c2)), transport remains quantized when calculated via the instantaneous soliton method. However, dynamically evolved soliton trajectories exhibit position jumps (Fig. 2(c2)) due to emergent self-intersecting bands in the intermediate nonlinear regime. This band topology, characterized by loop structures [85], is visualized in Fig. 2(c1).
We next examine how NNN tunneling amplitudes $t_{a}$ and $t_{b}$ in Hamiltonian (1) influence quantized soliton transport, maintaining fixed weak nonlinearity ($g=1$). These effects are systematically investigated through Fig. 3. We first analyze the case $t_{a}=0.4$, $t_{b}=-0.1$. As demonstrated in Fig. 3(a), gapless edge states persist due to preserved inversion symmetry, resulting in stable soliton oscillations (Fig. 3(b)). Crucially, the ground-state soliton exhibits a pumping value of $2$ at this parameter point—consistent with Chern number $\mathcal{C}=-1$ (corresponding to two atoms per unit cell)—and displays perfect adiabatic following, where dynamically evolved solitons precisely match their instantaneous counterparts. When the absolute values of $t_{b}$ becomes comparable to $t_{a}$ ($t_{a}=0.5$, $t_{b}=-0.5$), inversion symmetry breaks. This symmetry breaking opens a bandgap in the $\mu$-spectrum (Fig. 3(c)) and hybridizes edge states into bulk modes, thereby disrupting the adiabatic pathway as evidenced in Fig. 3(d). By selecting distinct NNN tunneling amplitudes $\{t_{a},t_{b}\}$ and superposing their corresponding soliton trajectories as shown in Fig. 3(e), we observe that all pumping pathways exhibit perfect adiabatic following except along the symmetry-broken line $t_{a}=-t_{b}=0.5$. Crucially, adiabatic breakdown occurs exclusively on this $t_{a}=-t_{b}$ trajectory. This reveals how NNN tunneling parameters reconfigure the system’s topological phase diagram.
Figure 4: NNN tunneling modulation of topological phases: (a,c) Nonlinear eigenvalue spectra ($\mu$); (b,d,e) Quantized soliton transport dynamics. Fixed parameters: $J=1$, $\delta=0.5$, $\Delta=1$, $g=6.215$, $T=2000\pi$, $\omega_{\text{d}}=10^{-3}$. Parameter configurations: (a,b) $t_{a}=0.4$, $t_{b}=-0.1$; (c,d) $t_{a}=0.5$, $t_{b}=-0.5$.